1. A force of 48.0 N is required to start a 6.0 Kg box moving across a horizontal
concrete floor.
a) What is the coefficient of static friction between box and floor?
b) If the 48.0 N force continues, the box accelerates at 0.70m/s2
, what is the
coefficient of kinetic friction?
1
Expert's answer
2020-05-21T13:40:51-0400

Given data :

Force "F = 48N"


Mass of box "m = 6 Kg"


Normal Reaction "N=mg= 6 \times 9.81 N = 58.86 N"


Frictional Force "F_{Frictiona} = F = 48 N"


Static Friction "\mu_{Static} = \frac {F_{Friction}} {N}"


Static Friction:


"\mu_{Static} = \frac {48} {58.86} = 0.815"


Kinetic Friction :


"\mu_{Kinetic} = \frac {F-F_{Acceleration}} {N}"


"\mu_{Kinetic} = \frac {48 - 6\times 0.70 } {58.86} = 0.744"